OpenStudy (anonymous):

Integral of dx/(3x^2+1)

OpenStudy (watchmath):

\(\int \frac{dx}{(\sqrt{3}x)^2+1)}=\frac{1}{\sqrt{3}}\arctan(\sqrt{3}x)+C\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

What is this question in relation to?